OpenStudy (anonymous):

60 km/hr what is that in m/s

OpenStudy (anonymous):

conversion trick 1 km/hr= 8/15 m/s i.e. 60 km/hr=8*60/15 m/s =32 m/s

OpenStudy (anonymous):

wat about doing it using metres and seconds like: km = 1000 m hrs = 60*60 = 3600 sec how would that work?

OpenStudy (anonymous):

say. 60 km/hr ======= /1000 * 3600 is this close to being right?

OpenStudy (anonymous):

almost just u missed its 1000/3600

OpenStudy (anonymous):

ok so its: 60*1000/3600 but yours was 32 m/s

OpenStudy (anonymous):

60*1000*3600 = 16.67 m/s

OpenStudy (mathstudent55):

\(60 \dfrac{km}{hr} \times \dfrac{1000 ~m}{1 ~km} \times \dfrac {1 ~hr}{60 ~min} \times \dfrac {1 ~min} {60 ~sec} \)